TREX: Logbook Entries

TREX: data-system Messages: 11 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
268 Mon 01-May-2006West data copiedwestpoulos
267 Mon 01-May-2006South data copiedsouthpoulos
250 Wed 26-Apr-2006Redundant copies for exthd7, 8, 9, 10allpoulos
196 Sat 15-Apr-2006Short local storage outagecentralmilitzer
188 Fri 14-Apr-2006Local storage retrievedwestmilitzer
170 Mon 10-Apr-2006restarted data systemsallmaclean
114 Mon 27-Mar-2006Copy local storage to aster over the networkwesthorst
108 Sun 26-Mar-2006dsm server downbasehorst
45 Sun 12-Mar-2006central died last nightcentraloncley
15 Tue 28-Feb-2006usb drive installsouthoncley
14 Tue 28-Feb-2006usb drive restartwestoncley

268: data-system, Site west, Mon 01-May-2006 16:09:27 PDT, West data copied
The download of the onboard pocketec data to exthd7 from West was started as of ~1610LT.

Based on data storage levels on exthd7 and exthd9 the data from Central tomorrow may need to be placed on exthd9.
267: data-system, Site south, Mon 01-May-2006 12:42:50 PDT, South data copied
The south pocketec data was copied to /media/exthd7 in the 'south' subdirectory. This is the final download of south data.
250: data-system, Site all, Wed 26-Apr-2006 16:53:43 PDT, Redundant copies for exthd7, 8, 9, 10
All hard-drives exthd7, 8, 9, 10 have been check_dirs and updated (through the 25 Apr hot-film run and the latest pocketec downloads from each tower).
196: data-system, Site central, Sat 15-Apr-2006 12:59:55 PDT, Short local storage outage
15Apr06, 11:15-12:06PDT

When I reconfigured power at central and booted the DSM back up, the
pocketec drive failed start up properly.
At 12:05 I cycled power on the pocketec and it started up again.

188: data-system, Site west, Fri 14-Apr-2006 10:51:58 PDT, Local storage retrieved
Measured electric fence battery = 12.66
Retrieved local storage on pocketec for archiving at base.  Swapped in clean drive.
Grabbed soil sample for Boulder tests.
170: data-system, Site all, Mon 10-Apr-2006 11:59:51 PDT, restarted data systems
Towers and base are now running new NIDS code

New code has the StatisticsProcessor and support for
prompting serial sensors (needed for SOAD charge monitor).

114: data-system, Site west, Mon 27-Mar-2006 14:13:29 PST, Copy local storage to aster over the network
Created a new directory on /mnt/exthd8/projects/TREX called local.
Downloaded data with, e.g.
scp root@west:/var/tmp/usb/TREX/west_20060326_000000.dat 
108: data-system, Site base, Sun 26-Mar-2006 09:47:08 PST, dsm server down
dsm server stopped yesterday at 0225 GMT, March 25; restarted this morning.

45: data-system, Site central, Sun 12-Mar-2006 16:51:17 PST, central died last night
Central died last night presumably due to power (we added the load of a 
second etherant plus the prometheus yesterday evening).  Came up by itself in
the morning -- probably about 1 hour lost.  Also rebooted during the day 
when new batteries (which hopefully will help the power problem) were installed
at about 1pm(?).  It appears that the local usb kept up with these cycles.

As usual, lost 15m sonic data until manually recycled both times.

15: data-system, Site south, Tue 28-Feb-2006 19:31:30 PST, usb drive install
The usb drive had never been installed at South.  Started it about 19:00 -- our
first data from South :(.

14: data-system, Site west, Tue 28-Feb-2006 19:29:43 PST, usb drive restart
Visited west site ~1800-1840 to restart the usb drive (which needed fsck).
Also repointed the parabolic antenna to South, since south beacon was visible.
Also checked guy wire tensions (see earlier log entry).